OpenStudy (anonymous):

Please and thank you for helping. Choose the equation of the line passing through the point (1, -5)and parallel to y = 2x + 4. Please show me how to get the answer and what the answer is I think it is 7=2x-3

OpenStudy (anonymous):

y = -2x + 3 y = 2x - 3 y = -2x + 7 y = 2x - 7

myininaya (myininaya):

What kinds of slopes do parallels line have?

OpenStudy (anonymous):

Geography is my thing not math so I dont know

OpenStudy (anonymous):

Well, if two lines are parallel, they have the same what?

OpenStudy (radar):

Parallel lines have ___________ slopes. Fill in the blalnk

OpenStudy (anonymous):

same idfk

myininaya (myininaya):

thats right! :)

myininaya (myininaya):

and we know a point on that line (1,-5) so use this formula \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

I dont have time to work it out I am on 16 out of 21 I have 2 minutes left and Im horrid at math.

myininaya (myininaya):

well you said that m is 2 and that x1 is 1 and x2 is -5 just plug them and then you have your line

OpenStudy (anonymous):

Im thankful ud help explain but I have to hurry I cant skip questions.

myininaya (myininaya):

you can put it in this form y=mx+b pretty easily after that

myininaya (myininaya):

ok